OpenStudy (anonymous):

find the vertex y-int x-int -x^2+6x-5

OpenStudy (asadkarim7):

set y = 0 to obtain x intercept set x=0 to obtain y intercept

OpenStudy (anonymous):

solve by factoring \[-x^2+6x-5=0\] \[x^2-6x+5=0\] \[(x-1)(x-5)=0\] \[x-1=0\] \[x=1\] \[x-5=0\] \[x=5\] your x - intercepts are 1 and 5

OpenStudy (anonymous):

whats the vertex

OpenStudy (anonymous):

magic formula is \[-\frac{b}{2a}=-\frac{6}{2\times -1}=3\]

OpenStudy (anonymous):

first coordinate of vertex is 3. second is what you get when you replace x by 3